# TiCoSS: Tightening the Coupling between Semantic Segmentation and Stereo Matching within A Joint Learning Framework

### v1

@@ IV-D1 Semantic Segmentation Performance
The qualitative experimental results on the KITTI and vKITTI2 datasets are presented in Figs. 2 and 3, respectively, while the quantitative experimental results on these two datasets are given in Tables I and II, respectively. Additionally, in Fig. 4, we present qualitative experimental results under various adverse weather conditions, including foggy, rainy, low-light, and sunny scenarios, to validate our model’s performance and robustness in complex situations.
These results suggest that TiCoSS outperforms all other SoTA single-modal and feature-fusion networks (including both CNN-based and Transformer-based methods) across all evaluation metrics on both datasets. Specifically, compared with S3M-Net, the SoTA joint learning method, TiCoSS demonstrates substantial improvements on the KITTI dataset, achieving increases of $9.68\%$ in mAcc, $10.57\%$ in mIoU, $2.71\%$ in fwIoU, and $1.20\%$ in mFSc, respectively. Similarly, on the vKITTI2 dataset, it outperforms other networks across all evaluation metrics, with improvements of $3.88\%$ in mAcc, $5.08\%$ in mIoU, $0.62\%$ in fwIoU, and $0.26\%$ in mFSc, respectively. Particularly, as observed in Figs. 2 and 3, TiCoSS achieves more accurate predictions on distant regions as well as object boundaries and is capable of providing more fine-grained semantic segmentation details compared to S3M-Net. In addition, as shown in Fig. 4, our TiCoSS also provides more accurate segmentation results in complex scenarios, such as the foggy and raining conditions. This demonstrates our model’s superior performance and robustness in complex scenarios, compared to other SoTA networks. We attribute these improvements to the tighter coupling between the two tasks within our joint learning framework, which effectively leverages the geometric information extracted from our predicted disparity maps. This enhances the integration of contextual and geometric features through our proposed TGF strategy, further improving the ability of handling heterogeneous features.
@@ IV-D2 Stereo Macthing Performance
The qualitative experimental results on the KITTI and vKITTI2 datasets are illustrated in Figs. 5 and 6, respectively, while the quantitative experimental results on these two datasets are provided in Table III. Since the primary focus of this study is to improve semantic segmentation performance, the stereo matching performance of TiCoSS is slightly better than that of S3M-Net. Specifically, compared to S3M-Net, TiCoSS demonstrates improvements of $3.64\%$ in EPE and $2.47\%$ in PEP 3.0 on the KITTI dataset. Additionally, on the vKITTI2 dataset, it achieves improvements of $5.26\%$ in EPE and $1.26\%$ in PEP 1.0. These improvements can be attributed to the tighter coupling between these two tasks, resulting in more comprehensive geometric features with informative contextual information compared to S3M-Net. Additionally, by minimizing the CT loss, our model focuses more on areas with inconsistent disparities and achieves improved performance in occluded regions, as depicted in Fig. 6.
@@ IV-E Ablation Studies
TABLE IV: Quantitative comparisons between our proposed TGF strategy and two SoTA feature fusion strategies, ASFF and GFF, on the KITTI 2015[30] dataset. “Baseline”: S3M-Net w/o SCG loss [16]. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@ IV-E1 Heterogeneous Feature Fusion Strategy
Extensive experiments are conducted on the KITTI 2015[30] dataset to validate the effectiveness of our proposed TGF strategy, compared with other two SoTA feature fusion strategies: Adaptive Spatial Feature Fusion (ASFF) [61] and GFF [52]. As presented in Table IV, TGF outperforms both ASFF and GFF, with an increase in mIoU by up to $7.93\%$, and an increase in fwIoU by $8.26\%$. These compelling results demonstrate the capability of TGF for effective feature extraction and for selective feature propagation. Additionally, we demonstrate the rationality of employing feature fusion strategies within each encoding branch. It is evident that all models perform best when feature fusion is adopted in both branches, demonstrating the necessity to selectively extract and fuse heterogeneous features within both encoding branches.
@@ IV-E2 Deep Supervision Strategy
We first compare the performance of our proposed HDS strategy with Single-resolution Deep Supervision (SDS) [39], Full-scale Deep Supervision (FDS) [53], and the combination of them (referred to SDS + FDS). As presented in Table V, our proposed HDS achieves the SoTA performance across both evaluation metrics, with improvements of up to $5.59\%$ in mIoU and $2.02\%$ in fwIoU. These compelling results can be attributed to the improved interactions among auxiliary classifiers, leveraging fused features with the richest local spatial details in the main branch to guide deep supervision across side branches. Additionally, it is noteworthy that the straightforward combination of SDS and FDS results in a marginal improvement compared to using FDS alone. Additionally, SDS is adopted only at the highest resolution, where features have a low number of channels, resulting in almost no additional memory usage. Therefore, we incorporate SDS into our HDS. Furthermore, we conduct an additional ablation study to evaluate the performance of HDS when utilizing guidance features in different layers and of various types. As shown in Table VI, the fused features at the shallowest layer results in the best performance compared to another two. We attribute this superior performance to the rich, fine-grained local spatial details $\mathbf{F}_{1}^{F}$ contains, which are essential for semantic segmentation.
@@ IV-E3 Loss Function
We systematically incorporate each element into the CT loss to assess its influence on the overall performance. Experimental results presented in Table VII validate the effectiveness of each component of our proposed CT loss. It is evident that the DIA and DSCC losses make significant contributions to semantic segmentation. Notably, when the entire joint learning framework is trained by minimizing the CT loss, TiCoSS achieves its best performance on the KITTI dataset, attaining a maximum mIoU of $63.91\%$. Additionally, to maximize the effectiveness of our proposed loss function, we first conduct an ablation study on the selection of loss weight $\alpha$ in (15). Fig. 7 (a) shows the mAcc and mIoU values with respect to different $\alpha$ within the range of $0.0$ to $2.0$. It can be obviously observed that when $\alpha=1.5$, TiCoSS achieves the best overall performance for both evaluation metrics. Following the selection of $\alpha$, we conduct another ablation study to determine $\beta$ in (16). Fig. 7 (b) demonstrates that $\mathbf{\beta=1.0}$ is the best choice. Further weight tuning is possible, but it should be approached cautiously, especially when dealing with limited data to avoid over-fitting.
@@ IV-E4 Combined use of A, B and C
We explore the rationality of each contribution adopted in our TiCoSS. As presented in Table VIII, we sequentially incorporate each novel component to quantify its impact on the overall performance. Please note that our CT loss is based on HDS strategy. Therefore, no ablation study needs to be conducted solely with the CT loss. The results demonstrate that each component employed in our model contributes to an improvement in the overall performance, and the network achieves its peak performance when all novel contributions (TGF, HDS, and CT loss) are leveraged, demonstrating the effectiveness of our design.
TABLE V: Ablation study on our HDS strategy on the KITTI 2015 [30] dataset. “Baseline”: S3M-Net [16] enhanced by our TGF strategy. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
TABLE VI: Ablation study on the selection of the guidance features employed in our HDS strategy on the KITTI 2015 [30] dataset. “Feature Layer”: the layer index of the guidance feature, with options of 1, 2, and 3. “GF”: Geometric Features, “CF”: Contextual Features, “FF”: Fused Features, obtained by performing an element-wise summation of the geometric and contextual features. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
TABLE VII: Ablation study to validate the effectiveness of the three semantic segmentation losses within our CT loss on the KITTI 2015 [30] dataset. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
63.91 86.82
TABLE VIII: Ablation study on the three contributions on the KITTI 2015 [30] dataset. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@ V Conclusion and Future Work
This article introduced TiCoSS, a novel joint learning framework designed to tighten the coupling between semantic segmentation and stereo matching tasks. We have made three significant contributions in this work: (1) an effective feature fusion strategy that complements the contextual features with useful geometric features through a series of selective inheritance gates, enabling a tightly-coupled duplex encoder, (2) a novel hierarchical deep supervision strategy that improves the interactions among auxiliary classifiers, and (3) a coupling tightening joint learning loss that focuses on further tightening the coupling of these two tasks at the output level. The effectiveness of each contribution was validated through extensive experiments on both virtual and real-world datasets. Comprehensive comparisons with other state-of-the-art algorithms unequivocally demonstrate the superiority of TiCoSS. Nevertheless, TiCoSS requires both semantic and disparity annotations, and collecting data with such ground truth remains a labor-intensive process. Thus, exploring potential solutions like semi-supervised or few-shot semantic segmentation, and un/self-supervised stereo matching is a promising avenue for future research. Besides, we believe that further optimization of model complexity is necessary before deploying TiCoSS in autonomous vehicles.

### v2

@@ IV-D1 Semantic Segmentation Performance
The qualitative experimental results on the KITTI and vKITTI2 datasets are presented in Figs. 2 and 3, respectively, while the quantitative experimental results on these two datasets are given in Tables I and II, respectively. We also provide additional quantitative experimental results evaluated by the mmsegmentation framework [61] in Table III. Additionally, in Fig. 4, we present qualitative experimental results under various adverse weather conditions, including foggy, rainy, low-light, and sunny scenarios, to validate the robustness of our model in complex situations.
These results suggest that TiCoSS outperforms all other SoTA single-modal and feature-fusion networks (including both CNN-based and Transformer-based methods) across all evaluation metrics on both datasets. Specifically, compared with S3M-Net, the SoTA joint learning method, TiCoSS demonstrates substantial improvements on the KITTI dataset, achieving increases of $9.68\%$ in mAcc, $10.57\%$ in mIoU, $2.71\%$ in fwIoU, and $1.20\%$ in mFSc, respectively. Similarly, on the vKITTI2 dataset, it outperforms other networks across all evaluation metrics, with improvements of $3.88\%$ in mAcc, $5.08\%$ in mIoU, $0.62\%$ in fwIoU, and $0.26\%$ in mFSc, respectively. Particularly, as observed in Figs. 2 and 3, TiCoSS achieves more accurate predictions on distant regions as well as object boundaries and is capable of providing more fine-grained semantic segmentation details compared to S3M-Net. In addition, as shown in Fig. 4, our TiCoSS also provides more accurate segmentation results in complex scenarios, such as the foggy and raining conditions. We attribute these improvements to the tighter coupling between the two tasks within our joint learning framework, which effectively leverages the informative geometric information extracted from our predicted disparity maps. This enhances the integration of contextual and geometric features through our proposed TGF strategy, further improving the ability of handling heterogeneous features.
@@ IV-D2 Stereo Macthing Performance
The qualitative experimental results on the KITTI and vKITTI2 datasets are illustrated in Figs. 5 and 6, respectively, while the quantitative experimental results on these two datasets are provided in Table IV. Since the primary focus of this study is to improve semantic segmentation performance, the stereo matching performance of TiCoSS is slightly better than that of S3M-Net. Specifically, compared to S3M-Net, TiCoSS demonstrates improvements of $3.64\%$ in EPE and $2.47\%$ in PEP 3.0 on the KITTI dataset. Additionally, on the vKITTI2 dataset, it achieves improvements of $5.26\%$ in EPE and $1.26\%$ in PEP 1.0. These improvements can be attributed to the tighter coupling between these two tasks, resulting in more comprehensive geometric features with informative contextual information compared to S3M-Net. Additionally, by minimizing the CT loss, our model focuses more on areas with inconsistent disparities and achieves improved performance in occluded regions, as depicted in Fig. 6.
@@ IV-E Ablation Studies
TABLE V: Quantitative comparisons between our proposed TGF strategy and two SoTA feature fusion strategies, ASFF and GFF, on the KITTI 2015[33] dataset. “Baseline”: S3M-Net w/o SCG loss [16]. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@ IV-E1 Heterogeneous Feature Fusion Strategy
Extensive experiments are conducted on the KITTI 2015[33] dataset to validate the effectiveness of our proposed TGF strategy, compared with other two SoTA feature fusion strategies: Adaptive Spatial Feature Fusion (ASFF) [65] and GFF [54]. As presented in Table V, TGF outperforms both ASFF and GFF, with an increase in mIoU by up to $7.93\%$, and an increase in fwIoU by $8.26\%$. These compelling results demonstrate the capability of TGF for effective feature extraction and for selective feature propagation. Additionally, we demonstrate the rationality of employing feature fusion strategies within each encoding branch. It is evident that all models perform best when feature fusion is adopted in both branches, demonstrating the necessity to selectively extract and fuse heterogeneous features within both encoding branches.
@@ IV-E2 Deep Supervision Strategy
We first compare the performance of our proposed HDS strategy with Single-resolution Deep Supervision (SDS) [42], Full-scale Deep Supervision (FDS) [55], and the combination of them (referred to SDS + FDS). As presented in Table VI, our proposed HDS achieves the SoTA performance across both evaluation metrics, with improvements of up to $5.59\%$ in mIoU and $2.02\%$ in fwIoU. These compelling results can be attributed to the improved interactions among auxiliary classifiers, leveraging fused features with the richest local spatial details in the main branch to guide deep supervision across side branches. Additionally, it is noteworthy that the straightforward combination of SDS and FDS results in a marginal improvement compared to using FDS alone. Additionally, SDS is adopted only at the highest resolution, where features have a low number of channels, resulting in almost no additional memory usage. Therefore, we incorporate SDS into our HDS. Furthermore, we conduct an additional ablation study to evaluate the performance of HDS when utilizing guidance features in different layers and of various types. As shown in Table VII, the fused features at the shallowest layer results in the best performance compared to another two. We attribute this superior performance to the rich, fine-grained local spatial details $\mathbf{F}_{1}^{F}$ contains, which are essential for semantic segmentation.
@@ IV-E3 Loss Function
We systematically incorporate each element into the CT loss to assess its influence on the overall performance. Experimental results presented in Table VIII validate the effectiveness of each component of our proposed CT loss. It is evident that the DIA and DSCC losses make significant contributions to semantic segmentation. Notably, when the entire joint learning framework is trained by minimizing the CT loss, TiCoSS achieves its best performance on the KITTI dataset, attaining a maximum mIoU of $63.63\%$. Additionally, to maximize the effectiveness of our proposed loss function, we first conduct an ablation study on the selection of loss weight $\alpha$ in (15). Fig. 7 (a) shows the mAcc and mIoU values with respect to different $\alpha$ within the range of $0.0$ to $2.0$. It can be obviously observed that when $\mathbf{\alpha=1.5}$, TiCoSS achieves the best overall performance for both evaluation metrics. Following the selection of $\alpha$, we conduct another ablation study to determine $\beta$ in (16). Fig. 7 (b) demonstrates that $\mathbf{\beta=1.0}$ is the best choice. Further weight tuning is possible, but it should be approached cautiously, especially when dealing with limited data to avoid over-fitting.
@@ IV-E4 Combined use of A, B and C
We explore the rationality of each contribution adopted in our TiCoSS. As presented in Table IX, we sequentially incorporate each novel component to quantify its impact on the overall performance. Please note that our CT loss is based on HDS strategy. Therefore, no ablation study needs to be conducted solely with the CT loss. The results demonstrate that each component employed in our model contributes to an improvement in the overall performance, and the network achieves its peak performance when all novel contributions (TGF, HDS, and CT loss) are leveraged, demonstrating the effectiveness of our design.
TABLE VI: Ablation study on our HDS strategy on the KITTI 2015 [33] dataset. “Baseline”: S3M-Net [16] enhanced by our TGF strategy. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
TABLE VII: Ablation study on the selection of the guidance features employed in our HDS strategy on the KITTI 2015 [33] dataset. “Feature Layer”: the layer index of the guidance feature, with options of 1, 2, and 3. “GF”: Geometric Features, “CF”: Contextual Features, “FF”: Fused Features, obtained by performing an element-wise summation of the geometric and contextual features. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
TABLE VIII: Ablation study to validate the effectiveness of the three semantic segmentation losses within our CT loss on the KITTI 2015 [33] dataset. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@
63.63 86.68
TABLE IX: Ablation study on the three contributions on the KITTI 2015 [33] dataset. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. The best results are shown in bold font.
@@ V Conclusion and Future Work
This article introduced TiCoSS, a novel, high-performing, state-of-the-art joint learning framework designed to tighten the coupling between semantic segmentation and stereo matching tasks. We made three key contributions in this work: (1) an effective feature fusion strategy and a tightly-coupled duplex encoder, leveraging the informative spatial information to enhance the semantic segmentation task, (2) a novel hierarchical deep supervision strategy that improves the interactions among all auxiliary classifiers, and (3) a coupling tightening joint learning loss that focuses on further tightening the coupling of these two tasks at the output level. The effectiveness of each contribution was validated through extensive experiments. Notably, our TiCoSS achieves a significant increase in mIoU of more than 9% compared to prior arts. Despite its superior performance over existing approaches, TiCoSS still requires both semantic and disparity annotations, and collecting data with such ground truth remains a labor-intensive process. Thus, exploring potential solutions like semi-supervised or few-shot semantic segmentation, and un/self-supervised stereo matching is a promising avenue for future research. Additionally, we believe that further optimization of model complexity is necessary before deploying TiCoSS in autonomous vehicles.